OpenStudy (anonymous):

check all that apply. if cot theta= 3/4 and the terminal point determined by theta is in quadrant 3, then: (a) tan theta= 4/3 (b) csc theta = -5/3 (c) sin theta= 3/5 (d) cos theta = -3/5

OpenStudy (anonymous):

@jdoe0001 can you please help

OpenStudy (jdoe0001):

\(\bf cot(\theta)=\cfrac{3}{4}\implies \cfrac{adjacent}{opposite}\implies \cfrac{a}{b}\\ \quad \\ \textit{using the pythagorean theorem }{\color{blue}{ c}}^2=a^2+b^2\implies {\color{blue}{ c}}=\pm \sqrt{a^2+b^2}\) the pythagorean theorem doesn't tell us if the root will be negative or positive, but we know that the angle is in Quadrant III, that is, cosine, or "x", is negative and sine, or "y", is negative as well thus we can say that \(\bf cot(\theta)=\cfrac{3}{4}\implies \cfrac{adjacent}{opposite}\implies \cfrac{a=-3}{b=-4}\)

OpenStudy (jdoe0001):

once you have all sides, that is, "a", "b", and "c", then check which ones are correct from the given options, that is, which ones apply

OpenStudy (jdoe0001):

I should point out that also the pythagorean theorem doesn't tell us if the root will be negative or positive, for "c" doesn't really matter, since the radius is always positive

OpenStudy (anonymous):

Thanks!

OpenStudy (cwrw238):



OpenStudy (anonymous):

so would the answer be B and D?

OpenStudy (anonymous):

@cwrw238

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

hmmm what did you for "c" anyway?

OpenStudy (anonymous):

???

OpenStudy (jdoe0001):

\(\bf cot(\theta)=\cfrac{3}{4}\implies \cfrac{adjacent}{opposite}\implies \cfrac{a}{b}\\ \quad \\ \textit{using the pythagorean theorem }{\color{blue}{ c}}^2=a^2+b^2\implies {\color{blue}{ c}}=\pm \sqrt{a^2+b^2}\)

OpenStudy (jdoe0001):

keep in mind both "a" and "b" are negative

OpenStudy (anonymous):

C = 5?

OpenStudy (anonymous):

-5

OpenStudy (jdoe0001):

yeap, so well, is "5" since "c" is just the radius, thus is never negative so would the answer be B and D? \(\Large \checkmark\)

OpenStudy (anonymous):

so i was right

OpenStudy (jdoe0001):

well, le'ts notice A option as well \(\bf tan(\theta)=\cfrac{b}{a}\implies \cfrac{-4}{-3}\implies \cfrac{4}{3}\)

OpenStudy (jdoe0001):

so, that one applies too

OpenStudy (anonymous):

oh ok, i should've seen that one. it's the simplest one lol :p thanks btw

OpenStudy (jdoe0001):

yw

OpenStudy (jdoe0001):

hmmm come to think B doesn't apply

OpenStudy (anonymous):

what do you mean?

OpenStudy (jdoe0001):

\(\bf csc(\theta)=\cfrac{c}{b}\implies \cfrac{5}{-4}\implies -\cfrac{5}{4}\)

OpenStudy (anonymous):

oh you're right

OpenStudy (anonymous):

@jdoe0001 can you help me with one more please

OpenStudy (jdoe0001):

ok.... would be easier if you repost anew, thus more exposure anyway, and we can also revise each other

OpenStudy (cwrw238):

the answer is a and d right?

OpenStudy (anonymous):

yes